# MoEUT: Mixture-of-Experts Universal Transformers

### v1

@@ 5 Discussion and Limitations
MoE methods for Transformer language models have seen many recent advances. It is worth noting that despite many works on MoEs for Transformers (see, e.g., [24, 25, 26, 27, 28]), many of them have only focused on applying MoE to the feedforward layers. Notable exceptions are mixture-of-attention [27] and SwitchHead [31] (Sec. 2.2) which focus on MoE self-attention layers. In addition, until recently, it has been considered challenging to make MoE-based language models competitive against the dense baseline in the parameter-matched setting (unlike FLOPs/MAC-matched settings). In our MoEUT, we use $\sigma$-MoE [28], an MoE design that has been shown to be competitive even in such a setting.
@@
Limitation/Implementation. Our current implementation of the MoE layers uses the Triton kernel released with $\sigma$-MoE [31] for both the attention and the MLP parts of the model. This implementation is known to be suboptimal [31]. Compared to the standard Transformer with FlashAttention [56], our MoEUT model trains 1.5–2x slower. We estimate that with a more optimal implementation, the training speed should be close to the dense model, while inference should run faster.
Massive scaling. Our experiments used a modest training regime. This led to good models and allowed us to make rigorous comparisons, but scaling to massive training regimes for MoEUT remains an important avenue for future research. Such experiments would inevitably require a very large compute cluster, but the costs could also be mitigated somewhat by work optimizing our CUDA kernel.
@@ 6 Conclusion
We present MoEUT, a novel Mixture-of-Expert-based Universal Transformer (UT) model that addresses the fundamental limitation of the standard UT in terms of parameter-compute efficiency. MoEUT combines the most advanced MoE techniques for Transformers with our novel layer grouping method and layernorm scheme which are both shown to be crucial for shared-layer models. Our MoEUT allows for training competitive UTs on parameter-dominated tasks such as language modeling, while being significantly less compute intensive than the baselines without layer sharing. We break this long standing limitation of UTs for the first time. Experimentally our model outperforms dense baselines from 44M to 1B parameter scale on C4, SlimPajama, peS2o, and The Stack datasets. Zero-shot experiments confirm that the performance of MoEUT holds on downstream tasks, including BLiMP, CBT, Lambada, HellaSwag, PIQA and ARC-E. We hope that this work could revive research interest in Universal Transformers at larger scales, and serve as a stepping stone for achieving the superior generalization properties of UTs (typically limited to synthetic problems for now) in real-world settings.
@@ A.2 Growing Residual Norm In Standard Transformers
Figure 12: The update magnitude of different layers in a 44M parameter Transformer on C4.
A.3 Hyperparameters
@@
A.4 The Effects of $d_{\text{expert}}$ and $K$
@@
Figure 13: Performance of our 244M MoEUT on a held-out subset of C4 with expert sizes ($d_{\text{expert}}$) in the MLP layer.
We also analyzed the performance of our MoEUT model with different numbers of active experts in the MLP layer. This varies the amount of compute spent in the layer, and the number of active channels. We show the results in Fig. 14. Increasing the number of active experts always improves performance, but the returns are diminishing. We chose $G=16$ for our experiments because of efficiency reasons.
Figure 14: Performance of our 244M MoEUT on a held-out subset of C4 with different number of active experts for the $\sigma$-MoE layer.
A.5 Additional Analysis
Expert specialization to tokens/layers. Now conversely to the “per-token expert selection diversity” analysis presented in the main text, we analyze whether the experts activated by a token are layer-specific for that specific token. First, we count the total number of experts used by each token and measure the proportion of them that is used in each layer. The results are shown in Fig. 15. Here, we order the tokens (x-axis) by their frequency in the validation set (the same ordering is used for all layers). The first 6000 tokens are shown. We can see that most of the experts are present in all layers for the most popular tokens. This can be seen from the high score in the first few columns. We can also observe that experts tend to be more layer-specific for the less popular tokens. In addition, the set of experts selected in the early layers is typically less diverse than for the rest of the layers: only a small fraction of the used experts are present there. This is consistent with the findings for Layer 1 in Fig. 11.
Figure 15: Fraction of experts used in a specific layer out of all experts used by a given token.
A.6 Compute Requirements
@@
Note that the report is generated from Weights and Biases logs. Because of this, the duration might include effects of restarts because of SLURM preemption and effects for occasional slowdowns of the network drive. Additionally, in a few instances the Weights and Biases log buffer overflowed, making it impossible to determine the number of GPUs used for the experiment. In this case we report “??” for the experiment. Note that we only report the resource usage of the final experiments here. We estimate that the total cost of the failed experiments and preliminary runs is at least an order of magnitude higher than this.

### v2

@@ 5 Discussion and Limitations
MoE methods for Transformer language models have seen many recent advances. It is worth noting that despite many works on MoEs for Transformers (see, e.g., [24, 25, 26, 27, 28]), many of them have only focused on applying MoE to the feedforward layers. Notable exceptions are mixture-of-attention [27] and SwitchHead [31] (Sec. 2.2), which focus on MoE self-attention layers. In addition, until recently, it has been considered challenging to make MoE-based language models competitive against the dense baseline in the parameter-matched setting (unlike FLOPs/MAC-matched settings). In MoEUT, we use $\sigma$-MoE [28], an MoE design that has been shown to be competitive even in such a setting.
@@
Limitation/Implementation.
Our current implementation of the MoE layers uses the Triton kernel released with $\sigma$-MoE [31] for both the attention and the MLP parts of the model. This implementation is known to be suboptimal [31]. Compared to the standard Transformer with FlashAttention [56], our MoEUT model trains 1.5–2x slower. We estimate that with a more optimal implementation, the training speed should be close to the dense model, while inference should run faster.
Massive scaling.
Our experiments used a modest training regime. This led to good models and allowed us to make rigorous comparisons, but scaling to massive training regimes for MoEUT remains an important avenue for future research. Such experiments would inevitably require a very large compute cluster, but the costs could also be mitigated somewhat by work optimizing our CUDA kernel.
@@ 6 Conclusion
We present MoEUT, a novel Mixture-of-Expert-based Universal Transformer (UT) model that addresses the fundamental limitation of the standard UT in terms of parameter–compute efficiency. MoEUT combines the most advanced MoE techniques for Transformers with our novel layer grouping method and layernorm scheme, which are both shown to be crucial for shared-layer models. Our MoEUT allows for training competitive UTs on parameter-dominated tasks such as language modeling, while being significantly less compute intensive than the baselines without layer sharing. We break this long standing limitation of UTs for the first time. Experimentally our model outperforms dense baselines from 44M to 1B parameter scale on C4, SlimPajama, peS2o, and The Stack datasets. Zero-shot experiments confirm that the performance of MoEUT holds on downstream tasks, including BLiMP, CBT, Lambada, HellaSwag, PIQA and ARC-E. We hope that this work helps revive research interest in Universal Transformers at larger scales, and serves as a stepping stone for achieving the superior generalization properties of UTs (typically limited to synthetic problems for now) in real-world settings.
@@ A.2 Growing Residual Norm In Standard Transformers
Figure 12: The update magnitude of different layers in a 44M parameter Transformer on C4. The norm of the updates grows throughout the layers to compensate for the residual growth (see Sec. 2.3 for more details).
A.3 Zero-Shot Downstream Performance of SUT-variants
In addition to evaluating the perplexity of different SUT variants in Fig. 8, we also show their zero-shot downstream performance in Tab. 2. It can be seen that MoEUT consistently outperforms both SUT and SUT without ACT.
Table 2: Zero-shot downstream performance and perplexity of different SUT variants compared to MoEUT and the unshared baseline. MoEUT outperforms both SUT variants.
C4 44M Baseline 18.97 21.9% 73.5% 81.3% 28.3% 59.9% 31.7% 49.4%
MoEUT
18.30
23.2%
78.2%
81.1%
29.2%
61.3%
33.5%
51.1%
SUT 40.50 1.2% 65.3% 51.1% 26.4% 57.8% 31.9% 39.0%
SUT w.o. ACT 21.51 18.1% 72.8% 66.3% 27.5% 59.1% 32.5% 46.0%
244M Baseline 13.40
33.1%
78.5%
86.0%
34.5% 64.9%
36.9%
55.6%
MoEUT
13.24
30.6%
79.7%
85.3%
35.7%
65.2%
36.4% 55.5%
SUT 20.05 20.5% 71.0% 68.5% 28.2% 60.1% 32.7% 46.8%
SUT w.o. ACT 14.58 27.8% 77.0% 75.9% 32.7% 63.2% 35.5% 52.0%
peS2o 44M Baseline 11.46
13.2%
66.5% 68.6%
28.5%
56.3%
32.0%
44.2%
MoEUT
11.09
13.1%
68.7%
69.6%
28.3% 55.1% 31.4%
44.4%
SUT 25.04 0.5% 59.2% 38.1% 26.2% 55.0% 31.1% 35.0%
SUT w.o. ACT 12.68 11.7% 66.5% 53.9% 28.0% 56.1% 31.5% 41.3%
244M Baseline 8.55 18.7% 72.8%
78.0%
30.4%
56.3% 35.0% 48.5%
MoEUT
8.52
19.4%
73.5%
77.4% 30.1% 56.3%
35.6%
48.7%
SUT 20.44 0.5% 60.9% 42.8% 26.7% 55.3% 33.0% 36.5%
SUT w.o. ACT 9.31 16.8% 71.9% 64.8% 28.8%
57.3%
34.8% 45.7%
A.4 Hyperparameters
@@
A.5 The Effects of $d_{\text{expert}}$ and $K$
@@
Figure 13: Performance of our 244M MoEUT on a held-out subset of C4 with expert sizes ($d_{\text{expert}}$) in the MLP layer. The smallest expert size performs the best.
We also analyzed the performance of our MoEUT model with different numbers of active experts in the MLP layer. This varies the amount of compute spent in the layer, and the number of active channels. We show the results in Fig. 14. Increasing the number of active experts always improves performance, but the returns diminish. We chose $G=16$ for our experiments because of efficiency reasons.
Figure 14: Performance of our 244M MoEUT on a held-out subset of C4 with different number of active experts for the $\sigma$-MoE layer. Increasing the number of experts always helps, but the returns diminish.
A.6 Additional Analysis
Expert specialization to tokens/layers. Now conversely to the “per-token expert selection diversity” analysis presented in the main text, we analyze whether the experts activated by a token are layer-specific for that specific token. For this, we count the number of unique experts used by each token and compute the corresponding proportion for each layer. The results are shown in Fig. 15. Here, we order the tokens (x-axis) by their frequency in the validation set (the same ordering is used for all layers). The first 6000 tokens are shown.
We observe that for the most frequent tokens (toward the left part of the plot), high scores near 1.0 are obtained in multiple layers for a given token; this means that (almost) all experts used by that specific token are used in multiple layers. In contrast, we observe that experts tend to be more layer-specific for the less popular tokens (toward the right part of the plot). In addition, the set of experts selected in the early layers is typically less diverse than for the rest of the layers: only a small fraction of the used experts are present there. This is consistent with the findings for Layer 1 in Fig. 10.
Figure 15: Proportion of experts used in a specific layer out of all unique experts used by a given token. On the x-axis, the tokens are ordered by decreasing frequency of occurrence. The less frequent a tokens is, the more layer-specific are the experts used by that token.
A.7 Wall Clock Time and Memory Comparison
To show the real-world resource usage of our models directly, we run a controlled experiment on identical hardware with our 244M parameter model and the corresponding baselines.
We measured the training iteration time and memory usage on 8 V100 32GB GPUs. Here one “iteration” corresponds to an effective batch size of 64x1024 tokens for all models. The training iteration time was measured by using a batch size for each model that fits GPUs; models require either 1 or 2 gradient accumulation steps to achieve the effective batch size depending on their memory requirement. We measured the training time right after initialization and a warmup period. The memory usage is measured using 2 grad accumulation steps for all models for a fair comparison. Note that around  3GB of memory is used by the model parameters and optimizer state on each GPU. We show the results in Tab. 5.
Even though our MoEUT with the current kernel implementation is slower than the corresponding dense non-shared layer transformer, it is significantly faster and uses much less memory compared to alternative UT variants.
Table 5: Wall-Clock Time for the forward-backward pass and the total memory usage of our training loop with different 244M parameter models on 8 V100 GPUs, which a batch size of 64x1024 tokens. MoEUT is 1.7x slower with our suboptimal MoE kernel implementation than the standard transformer, but it is much faster than the other UT variants. It also uses much less memory, allowing training on larger scales.
Model ms/batch Memory usage
Non-shared Transformer 443 9.2 G
Naive UT 3559 25.9 G
MoEUT 772 9.0 G
SUT 1344 23.4 G
A.8 Compute Requirements
@@
Note that the report is generated from Weights and Biases logs. Because of this, the duration might include effects of restarts because of SLURM preemption and effects for occasional slowdowns of the network drive. Additionally, in a few instances the Weights and Biases log buffer overflowed, making it impossible to determine the number of GPUs used for the experiment. In this case, we report “??” for the corresponding experiment. Note that we only report the resource usage of the final experiments here. We estimate that the total cost of the failed experiments and preliminary runs is at least an order of magnitude higher than this.